### v2

# Decoding convolutional codes over finite rings. A linear dynamical systems approach

Abstract.
Observable convolutional codes defined over $\operatorname{\mathbb{Z}_{p^{r}}}$ with the Predictable Degree Property admits minimal input/state/output representations that behaves well under restriction of scalars. We make use of this fact to present Rosenthal’s decoding algorithm for these convolutional codes. When combined with the Greferath-Vellbinger algorithm and a modified version of the Torrecillas-Lobillo-Navarro algorithm, the decoding problem of convolutional codes over $\operatorname{\mathbb{Z}_{p^{r}}}$ reduces to selecting two decoding algorithms for linear block codes over a field. Finally, we analyze both the theoretical and practical error-correction capabilities of the combined algorithm.
Key words and phrases:
codes over rings; dynamical systems; decoding
2020 Mathematics Subject Classification:
94B10; 94B35; 93B20;
This work is part of the project TED2021-131158A-I00, funded by MCIN/ AEI/10.13039/501100011033 and by the European Union “NextGenerationEU”/PRTR
1. Introduction
Peter Elias introduced convolutional codes in 1955 [1], and Forney gave the first algebraic theoretical approach to convolutional codes in [3]. Within the theory of convolutional codes, two fundamental problems arise: the construction of convolutional codes with good observability and distance properties, and the decoding problem. The first problem concerns designing convolutional codes that maximize their error-correction capabilities while maintaining efficient encoding and decoding properties. The second problem, the decoding, focuses on retrieving the original information sent through a channel, even in the presence of errors caused by the transmission process. Addressing both problems is crucial to improving the reliability and efficiency of convolutional communication systems.
One way to face both problems is to approach the theory of convolutional codes through other mathematical objects. One of the most studied approaches is the relation of convolutional codes and linear dynamical systems by the well-known input/state/output (I/S/O) representations, introduced and developed in [21, 22]. The existence of I/S/O representations of a convolutional code enables the construction of algorithms for the decoding process by exploiting the algebraic dependency between the information vector and the output vector solving systems of linear equations. If the transmission is over a noisy channel, the first decoding algorithm based on I/S/O representations was given in [23]. This algorithm retrieves the original information encoded with errors and is computationally more efficient than the Viterbi algorithm for certain families of convolutional codes. Based on this work, the decoding problem of convolutional codes over noisy channels using I/S/O representations has been and continues to be extensively studied ([4, 12, 18]). For convolutional codes over an erasure channel, a decoding algorithm based on I/S/O representations, following the idea of [23], was proposed in [25, 24] and it was improved in [10]. The latter proposal has the advantage of reducing both the decoding delay and the computational effort required in the process.
In this article, we will focus on the decoding process over noisy channels. We studied how minimal I/S/O representations can be used to decode convolutional codes over finite rings over noisy channels. In particular, for convolutional codes over $\operatorname{\mathbb{Z}_{p^{r}}}$. The motivation for considering convolutional codes over finite rings lies in their potential applications in communication systems. In particular, finite rings, such as $\operatorname{\mathbb{Z}_{p^{r}}}$, provide a natural framework for modeling coding schemes in scenarios where additional algebraic structure is beneficial. Massey and Mittelholzer gave the first approach to convolutional codes over rings in [14] and [13]. Since then, there has been a large amount of literature on convolutional codes over rings ([2, 7, 9, 8, 15, 19], among others). We want to explore whether Rosenthal’s decoding algorithm for convolutional codes over finite fields can be extended over finite rings, specifically over $\operatorname{\mathbb{Z}_{p^{r}}}$. With this goal, we define convolutional codes as submodules of $\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ and recall the necessary conditions (the Predictable Degree Property (PDP)) for these codes to admit a minimal I/S/O representation ([17]). Then, we reduce the decoding problem of convolutional codes to the decoding of two linear block codes over $\operatorname{\mathbb{Z}_{p^{r}}}$ as in Rosenthal’s decoding algorithm for observable convolutional codes defined over a field. The paper also includes a performance evaluation of the algorithm, providing insights into the advantages and limitations of using convolutional codes over finite rings in practical applications.
The article is organized as follows. Section II includes an overview of convolutional codes and their relation with linear dynamical systems over $\operatorname{\mathbb{Z}_{p^{r}}}$. The main results are developed in Section III, and they are completed in Section IV with the study of the performance of the proposed decoding algorithm. Finally, the conclusions and the references are given.
2. Convolutional codes and I/S/O representations over $\operatorname{\mathbb{Z}_{p^{r}}}$
We fix a prime number $p\geq 2$ and a natural number $r\geq 1$. We denote by $\operatorname{\mathbb{Z}_{p^{r}}}$ the quotient ring $\mathbb{Z}/{p^{r}\mathbb{Z}}$ and by $\operatorname{\mathbb{Z}_{p^{r}}}[z]$ the ring of polynomials with coefficients in $\operatorname{\mathbb{Z}_{p^{r}}}$.
2.1. Convolutional Codes
In this section, we include some fundamental results about convolutional codes.
Let $k$ and $n$ be natural numbers such that $k\leq n$. An $(n,k)$ convolutional code over $\operatorname{\mathbb{Z}_{p^{r}}}$ is defined as a submodule
$\mathcal{C}\subset\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}$
of rank $k$ [17, Definition 1]. Elements of $\mathcal{C}$ are called codewords.
For an $(n,k)$ convolutional code $\mathcal{C}$ over $\operatorname{\mathbb{Z}_{p^{r}}}$, a convolutional encoder is a full rank (i.e. injective) matrix
$G(z)\in\text{Mat}_{n\times k}(\operatorname{\mathbb{Z}_{p^{r}}}[z])$
such that the column space of $G(z)$ is equal to $\mathcal{C}$.
On the other hand, two convolutional codes $\mathcal{C},\mathcal{C}^{\prime}\subset\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}$ are equivalent if there is a permutation $\sigma\in\mathfrak{S}_{n}$ acting on the codeword components such that $\sigma(\mathcal{C})=\mathcal{C}^{\prime}$ [17, Definition 12]. Two convolutional encoders $G(z),G^{\prime}(z)\in\text{Mat}_{n\times k}(\operatorname{\mathbb{Z}_{p^{r}}}[z])$ are equivalent if they determine the same convolutional code, i.e., if and only if there is an invertible matrix $U(z)\in\text{Mat}_{k\times k}(\operatorname{\mathbb{Z}_{p^{r}}}[z])$ such that $G(z)=G^{\prime}(z)U(z)$ [7, Theorem 1].
A desired property of a convolutional code is that not be non-catastrophic. An $(n,k)$ convolutional code $\mathcal{C}\subset\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}$ is said to be non-catastrophic or observable if the natural quotient map
$\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}\twoheadrightarrow\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}/\mathcal{C}$
has a right inverse, which means that $\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}/\mathcal{C}$ is free of rank $n-k$ [17, Remark 1] and [2, §IIIA]. In fact $\mathcal{C}$ is non-catastrophic if and only if one has the decomposition $\operatorname{\mathbb{Z}_{p^{r}}}[z]\cong\mathcal{C}\oplus\left(\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}/\mathcal{C}\right)$. In particular, if $\mathcal{C}$ is non-catastrophic or observable then it is a free module. From now on, every convolutional code will be assumed to be a free module.
The $i$-th constraint length, $\nu_{i}$, of a convolutional encoder $G(z)$ is defined as the maximum degree of the entries in the $i$-th column of $G(z)$. Without loss of generality, we may assume that $\nu_{1}\geq\dots\geq\nu_{k}$ [17, Definition 5]. The maximum constraint length, $\nu_{1}$, is called the memory of the encoder. The complexity (or degree) of $G(z)$ is given by $\delta_{G(z)}:=\sum_{i=1}^{k}\nu_{i}$. The complexity of $\mathcal{C}$, denoted by $\delta_{\mathcal{C}}$, is defined as
$\delta_{\mathcal{C}}:=\textrm{min}\{\delta_{G(z)}|\ G(z)\textrm{ encoder of }\mathcal{C}\}$
Clearly,
$\delta_{\mathcal{C}}\geq\textrm{max}\left\{\begin{array}[]{l|l}\textrm{maximumdegree}&\\ \textrm{of the full-size}&G(z)\textrm{ encoder of }\mathcal{C}\\ \textrm{minors of }G(z)&\end{array}\right\}=:\delta^{\prime}_{\mathcal{C}}$
As we mentioned in Introduction, a needed property in order to obtain minimal I/S/O representations of a convolutional code over $\operatorname{\mathbb{Z}_{p^{r}}}$ is the PDP. We recall the definition. Given a convolutional encoder $G(z)$ of an $(n,k)$ convolutional code $\mathcal{C}$ over $\operatorname{\mathbb{Z}_{p^{r}}}$, the following holds:
$\text{deg}(G(z)u(z))\leq\max\{\text{deg}(u_{i}(z))+\text{deg}(g_{i}(z))\},$
where the degree of a polynomial vector is the maximum degree among its components, and $g_{i}(z)$ refers to the $i$-th column of $G(z)$. The encoder $G(z)\in\text{Mat}_{n\times k}(\operatorname{\mathbb{Z}_{p^{r}}}[z])$ has the Predictable Degree Property (PDP) if the above inequality becomes an equality for every $u(z)$. The PDP is characterized by the condition that the matrix of column-wise maximum degree coefficients, $G_{h}$, is injective [17, Theorem 1]. It follows easily that if $\mathcal{C}$ is a convolutional code $\mathcal{C}$ over $\operatorname{\mathbb{Z}_{p^{r}}}$ that has a convolutional encoder $G(z)$ with the PDP then $\delta_{G(z)}=\delta_{\mathcal{C}}=\delta^{\prime}_{\mathcal{C}}.$ We say that a convolutional code is PDP if it admits an encoder with the PDP.
Finally, the free distance of a convolutional code $\mathcal{C}\subset\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}$ is defined as
$d_{\textrm{free}}(\mathcal{C})=\textrm{min}\{\omega_{H}(v(z))|\ v(z)\in\mathcal{C},v(z)\neq 0\}.$
where for $v(z)=\sum_{i\in\mathbb{N}_{0}}v_{i}z^{i},\omega_{H}(v(z))=\sum_{i\in\mathbb{N}_{0}}\omega_{H}(v_{i})$, $\omega_{H}(v_{i})$ being the Hamming weight of $w_{i}$ [16, §1].
2.2. I/S/O Representations
[…]
$\begin{split}\mathcal{C}=\{v(z)\in\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}:&\ \exists x(z)\in\operatorname{\mathbb{Z}_{p^{r}}}[z]^{\delta}\text{ such that }\\ &zKx(z)+Lx(z)+Mv(z)=0\},\end{split}$ (2.1)
and such that
  1. (1)
    $K$ is injective.
  2. (2)
    $(K,M)$ is surjective.
  3. (3)
    $(zK+L,M)$ is surjective.
This triple is called a minimal first-order representation of the convolutional code $\mathcal{C}$ and it is unique up to similarities [17]. Moreover, there is always a permutation $\sigma\in\mathfrak{S}_{n}$ such that any minimal first-order representation $(K,L,M)$ of the code $\sigma(\mathcal{C})$ is similar to a minimal first-order representation of the form
$\left(\left(\begin{array}[]{c}-I\\ 0\end{array}\right),\left(\begin{array}[]{c}A\\ C\end{array}\right),\left(\begin{array}[]{cc}0&B\\ -I&D\end{array}\right)\right).$ (2.2)
[…]
$\Sigma:=\left\{\begin{array}[]{rl}x_{t+1}&=Ax_{t}+Bu_{t},\\ y_{t}&=Cx_{t}+Du_{t},\\ v_{t}&=\left(\begin{array}[]{c}y_{t}\\ u_{t}\end{array}\right),\ x_{0}=0,\ x_{\gamma+1}=0.\end{array}\right.$ (2.3)
Again, a minimal I/S/O representation of an observable convolutional code with PDP is unique up to similarities. Note that this construction works in the other way around. Starting from a reachable and observable linear system $\Sigma=(A,B,C,D)$ over the ring, the associated convolutional code $\mathcal{C}$ obtained by $Ker(zK+L\mid M)$ is observable [17, Theorem 5].
Associated to a minimal I/S/O representation of an observable convolutional code with PDP there are two important scalar matrices:
  1. (1)
    The reachability matrix:
    $\Phi_{l}(\Sigma):=\left(A^{l-1}B,\ldots,AB,B\right),$
    which is surjective in case $l=\delta$. That is equivalent to the dynamical system $\Sigma=(A,B,C,D)$ being reachable.
  2. (2)
    The observability matrix:
    $\Psi_{l}(\Sigma):=\left(\begin{array}[]{c}C\\ CA\\ \vdots\\ CA^{l-1}\end{array}\right),$
    which is injective in case $l=\delta$. That is equivalent to the dynamical system $\Sigma=(A,B,C,D)$ being observable.
Remark 2.4.
For any pair of values $T,\Theta$ making $\Phi_{T}(\Sigma)$ and $\Psi_{\Theta}(\Sigma)$ surjective and injective, respectively, we may regard these matrices as parity-check and generator matrices, respectively, of the corresponding linear block codes over $\operatorname{\mathbb{Z}_{p^{r}}}$, $\textrm{Ker}(\Phi_{T}(\Sigma)),\ \textrm{Im}(\Psi_{\Theta}(\Sigma))$. Rosenthal’s algorithm [23] works under the assumption of the existence of such values $T,\Theta$ and reduces the decoding process of the convolutional code $\mathcal{C}$ to the decoding of the linear block codes $\textrm{Ker}(\Phi_{T}(\Sigma)),\ \textrm{Im}(\Psi_{\Theta}(\Sigma))$ defined over $\operatorname{\mathbb{Z}_{p^{r}}}$. Therefore, linear block codes over $\operatorname{\mathbb{Z}_{p^{r}}}$ and their decoding algorithms will play an important role in Rosenthal’s decoding procedure.
3. The decoding algorithm
Given an $(n,k)$ convolutional code $\mathcal{C}\subset\operatorname{\mathbb{Z}_{p^{r}}[z]}^{n}$ and a received word $\widehat{v}(z)\in\operatorname{\mathbb{Z}_{p^{r}}[z]}^{n}$, the decoding problem consists of finding the codewords $v(z)\in\mathcal{C}$ such that $\omega_{H}(v(z)-\widehat{v}(z))$ attaches the minimum
$\textrm{min}_{0\neq v(z)\in\mathcal{C}}\{\omega_{H}(v(z)-\widehat{v}(z))\}.$ (3.1)
If the free distance of $\mathcal{C}$ is $d$, and the error vector $e(z)\in\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}$ corresponding to $\widehat{v}(z)$ has Hamming weight less than or equal to $\lfloor\dfrac{d-1}{2}\rfloor$, then there is a unique codeword in the set defined in Eq. (3.1). This must be the codeword sent (this follows by general principles as in [20, Proposition 2.5]). In this case, we say that the decoding is complete.
Note that the codewords $v(z)\in\mathcal{C}$ contain the information of the whole message that has been sent. Thus, decoding with convolutional codes is performed once the whole message has been received. This makes a subtle difference, at least conceptually, with respect to the decoding process with linear block codes $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$. In this last case, the message is encoded block by block (independently) and transmitted block by block, so the decoding can be performed also block by block finding
$\textrm{min}_{0\neq v\in\mathcal{C}}\{\omega_{H}(v-\widehat{v})\}$
without having the entire message.
Rosenthal’s decoding algorithm for convolutional codes over fields rests on the decoding of two linear block codes [23]. This implies that using this procedure we may decode by blocks of certain lengths once the whole message has been received as long as the weight of the error vector of each block is not too large. This will also be the situation in the case of convolutional codes over $\operatorname{\mathbb{Z}_{p^{r}}}$.
In this article, we propose to combine Rosenthal’s decoding algorithm with the Greferath-Vellbinger algorithm [6] and a modified version of the Torrecillas-Lobillo-Navarro algorithm [5].
First, we introduce some notation for the sake of clarity, and some preliminaries on linear block codes over $\operatorname{\mathbb{Z}_{p^{r}}}$.
Notation
The residue field of the finite ring $\operatorname{\mathbb{Z}_{p^{r}}}$ is the finite field $\mathbb{F}_{p}$.
Given an element $z\in\operatorname{\mathbb{Z}_{p^{r}}}$ with $p$-adic expansion $z=\sum_{j=0}^{r-1}z_{j}p^{j}$, we define $z^{(i)}=\sum_{j=0}^{i}z_{j}p^{j}$ and ${}^{(i)}z=\sum_{j=i}^{r-1}z_{j}p^{j}$. We extend these definitions to vectors and matrices with coefficients in $\operatorname{\mathbb{Z}_{p^{r}}}$. Finally, if an element $z\in\operatorname{\mathbb{Z}_{p^{r}}}$ can be written as $z=z^{\prime}p^{r-1}$ with $z^{\prime}\in\mathbb{F}_{p}$, we will use the notation $z^{\prime}=\dfrac{z}{p^{r-1}}$.
Finally, given an ordered set $c=(c_{0},\ldots,c_{l-1})$ and natural numbers $i<j\leq l-1$, we will denote by $c[i,j]$ the ordered set $(c_{i},\ldots,c_{j})$.
3.1. Linear block codes over $\operatorname{\mathbb{Z}_{p^{r}}}$
An $(n,k)$ linear block code over $\operatorname{\mathbb{Z}_{p^{r}}}$ is a submodule $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ of rank $k$. Linear block codes over $\operatorname{\mathbb{Z}_{p^{r}}}$ are not necessary free submodules. Thus, we say that the code $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ is free if it is a free submodule. On the other hand, the linear block code $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ is said to be splitting if it is a direct summand of $\operatorname{\mathbb{Z}^{n}_{p^{r}}}$. An encoder $G$ for the linear block code $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ is a full rank (injective) generator matrix $G$. These last two notions are in fact equivalent:
Lemma 3.2.
An $(n,k)$ linear block code over $\operatorname{\mathbb{Z}_{p^{r}}}$ admits an encoder if and only if it is splitting.
Proof.
It is sufficient to prove that if $\mathcal{C}$ admits an encoder then $\mathcal{C}$ is splitting.
Since $G$ is an full rank generator matrix it follows that it is injective and, therefore, its determinantal ideals are $\mathcal{U}_{\gamma}(G)=<1>$ for $\gamma\leq k$ and $\mathcal{U}_{\gamma}(G)=0$ for $\gamma>k$. This is equivalent [11, Theorem IV.32] to saying that quotient module $\operatorname{\mathbb{Z}^{n}_{p^{r}}}/\mathcal{C}$ is projective and hence free because $\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ is local. Therefore, we have $\operatorname{\mathbb{Z}^{n}_{p^{r}}}\cong\mathcal{C}\oplus\left(\operatorname{\mathbb{Z}^{n}_{p^{r}}}/\mathcal{C}\right)$ and the code $\mathcal{C}$ is splitting. ∎
The (minimum) distance of a linear block code $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ is defined as $d_{\textrm{min}}:=\textrm{min}_{0\neq c\in\mathcal{C}}\{\omega_{H}(c)\},$ where $\omega_{H}(c)$ is the Hamming weight of $c\in\mathcal{C}$. As for any block code, $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ can correct all weight errors up to $t$ if and only if $d_{\textrm{min}}=2t+1$. An important feature of any free (equivalently, splitting) linear block code $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$, is that the minimum distance coincides with the minimum distance of its restriction modulo $p$ ,
$d_{\textrm{min}}(\mathcal{C})=d_{\textrm{min}}(\mathcal{C}\textrm{ mod }p).$
Given an $(n,k)$ free linear block code over $\operatorname{\mathbb{Z}_{p^{r}}}$ of rank $k$, $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$, a parity-check matrix $H\in\text{Mat}_{n-k\times n}(\operatorname{\mathbb{Z}_{p^{r}}})$ and a vector $e\in\operatorname{\mathbb{Z}^{n}_{p^{r}}}$, its syndrome is $He\in\mathbb{Z}^{n-k}_{p^{r}}$ while its coset is $e+\mathcal{C}\in\operatorname{\mathbb{Z}^{n}_{p^{r}}}$. The (Hamming) weight of the coset $e+\mathcal{C}$ is $\omega_{H}(e+\mathcal{C}):=\textrm{min}_{e^{\prime}\in e+\mathcal{C}}\{\omega_{H}(e^{\prime})\}$. A coset leader of $e+\mathcal{C}$ is an element $e^{\prime}\in e+\mathcal{C}$ such that $\omega_{H}(e^{\prime})=\omega_{H}(e+\mathcal{C})$. If $e\in\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ has weight $\omega_{H}(e)\leq\lfloor\dfrac{d-1}{2}\rfloor$ then $e$ is the unique coset leader in its coset.
3.2. Greferath-Vellbinger (GV) algorithm for free linear block codes over $\operatorname{\mathbb{Z}_{p^{r}}}$
Let $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ be an $(n,k)$ free linear block code and $G\in\textrm{Mat}_{n\times k}(\operatorname{\mathbb{Z}_{p^{r}}})$ an encoder for $\mathcal{C}$. Suppose that the transmitted information word is $u\in\operatorname{\mathbb{Z}_{p^{r}}^{k}}$, the codeword is $v=Gu\in\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ and that the received codeword is $\widehat{v}\in\operatorname{\mathbb{Z}^{n}_{p^{r}}}$. Let $e:=\widehat{v}-v$ be the error vector. Assume we have a decoding algorithm $\partial$ for the linear code $\mathcal{C}_{0}:=\mathcal{C}\ \textrm{mod}(p)$. The GV algorithm computes the coefficients of the $p$-adic expansions of $u,e$ iteratively as stated in Algorithm 1 (see [6, pp. 1290]). It corrects any error pattern $e$ for which each component of its $p$-adic expansion is correctable by $\partial$.
Algorithm 1 GV algorithm
Parameters: $G\in\textrm{Mat}_{n\times k}(\operatorname{\mathbb{Z}_{p^{r}}})$ encoder of an $(n,k)$ free linear block code and a decoding algorithm $\partial$ for the linear code given by $G_{0}:=G\ \textrm{mod}(p)$.
Input: A received word $\widehat{v}=Gu+e\in\operatorname{\mathbb{Z}^{n}_{p^{r}}}$.
Output: $u,v,e\in\operatorname{\mathbb{Z}^{n}_{p^{r}}}$
$i=0$
while $i\leq k$ do
     $\delta_{i}\leftarrow\dfrac{\widehat{v}^{(i)}-G^{(i)}u^{(i-1)}-e^{(i-1)}}{p^{i}}$
     Solve $\delta_{i}=G_{0}u_{i}+e_{i}$ with $\partial$
     $i=i+1$
end while
return $e,u,v=Gu$
3.3. Torrecillas-Lobillo-Navarro (TLN) algorithm adapted to free linear block codes over $\operatorname{\mathbb{Z}_{p^{r}}}$
Let $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ be an $(n,k)$ linear block code and $H\in\textrm{Mat}_{q\times n}(\operatorname{\mathbb{Z}_{p^{r}}})$ a parity-check matrix for $\mathcal{C}$. Suppose that the transmitted codeword is $v\in\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ and that the received codeword is $\widehat{v}\in\operatorname{\mathbb{Z}^{n}_{p^{r}}}$. Let $e:=\widehat{v}-v$ be the error vector. Let $s:=H\widehat{v}=He$ be the syndrome of the received word. The TLN algorithm computes the coefficients of the $p$-adic expansion of the error $e$ iteratively. It assumes that the parity-check matrix $H$ is in standard form, meaning that there are matrices ${}^{(i)}\widehat{H}\in\textrm{Mat}_{q_{i}\times n}(\operatorname{\mathbb{Z}_{p^{r}}})$, $i=1,\ldots,r-1$ with $p$-adic expansion ${}^{(i)}\widehat{H}=\sum_{j=i}^{r-1}\Lambda_{ij}p^{j}$ such that $H=(^{(0)}\widehat{H}|^{(1)}\widehat{H}|\cdots|^{(r-1)}\widehat{H})$ (thus $\sum q_{i}=q$), and satisfying $\Lambda_{ii}$ is full rank for each $i$. Further, the algorithm assumes $r$ decoding algorithms are fixed, one for each $\Lambda_{ii}$.
Note that condition $\Lambda_{ii}$ is full rank for each $i$ makes ${}^{(i)}\widehat{H}\neq 0$ for each $i$ so that $q$ must satisfy $q\geq r$. In our case, we will have a full rank parity-check matrix $H$. We will show now how to get advantage of this fact to apply TLN algorithm without computing the standard form and using only one decoding algorithm for the code defined by $H_{0}:=H\textrm{mod}(p)$.
Let $H\in\textrm{Mat}_{n-k\times n}(\operatorname{\mathbb{Z}_{p^{r}}})$ be a full rank parity-check matrix of an $(n,k)$ free linear block code $\mathcal{C}\subset\operatorname{\mathbb{Z}^{n}_{p^{r}}}$ with $p$-adic expansion $H=\sum_{i=0}^{r-1}H_{i}p^{i}$. For each $i=0,\ldots,r-1$, we construct the matrices ${}^{(i)}\widehat{H}=p^{i}H$ and $\widehat{H}=(^{(0)}\widehat{H}|^{(1)}\widehat{H}|\cdots|^{(r-1)}\widehat{H})$. Clearly, ${}^{(0)}\widehat{H}=H$. Consider the $p$-adic expansions of these matrices, ${}^{(i)}\widehat{H}=\sum_{j=i}^{r-1}\Lambda_{ij}p^{j}$. The following facts are easy to check:
  1. (1)
    $\Lambda_{00}=H_{0}$ is full rank.
  2. (2)
    $\Lambda_{ij}=\Lambda_{0j-i}=H_{j-i}$. So $\Lambda_{ii}=H_{0}$ is full rank.
  3. (3)
    $\textrm{Ker}(\widehat{H})=\textrm{Ker}(H)=\mathcal{C}$.
Under these conditions, TLN algorithm can be applied to compute the error $e$ from the equation $\widehat{s}=\widehat{H}e$, where $\widehat{s}=(^{(0)}\widehat{s}|^{(1)}\widehat{s}|\cdots|^{(r-1)}\widehat{s})$ and ${}^{(i)}\widehat{s}:=p^{i}s$.
Let $e=\sum_{j=0}^{r-1}e_{j}p^{j}$ and $s=\sum_{j=0}^{r-1}s_{j}p^{j}$ be the $p$-adic expansions of $e$ and $s$, respectively. Note that the $p$-adic expansion of ${}^{(i)}\widehat{s}$ is ${}^{(i)}\widehat{s}=\sum_{j=i}^{r-1}s_{j-i}p^{j}$. Now, the equation ${}^{(i)}\widehat{s}=\widehat{H}e^{(i)}$ leads to the following equation (see [5, Eq. (12)] and condition 2) above)
$\sum_{j=i}^{r-1}s_{j-i}p^{j}=\sum_{j=i}^{r-1}\sum_{l=0}^{j-i}H_{j-i-l}e_{l}p^{j}$ (3.3)
Assume that $e_{0},\ldots,e_{l-1}$ are known. Then, from Eq. (3.3) it follows that there exist $\delta_{l}\in\mathbb{F}_{p}$ such that
$\begin{split}\Xi(l)&:=\big{(}\sum_{j=r-1-l}^{r-1}s_{j-r+1+i}p^{j}\big{)}\\ &-\big{(}\sum_{j=r-1-l}^{r-2}\sum_{i=0}^{j-r+1+l}H_{j-i-r+1+l}e_{i}p^{j}\big{)}\\ &-\big{(}\sum_{i=0}^{l-1}H_{l-i}e_{i}p^{r-1}\big{)}=\delta_{l}p^{r-1}\end{split}$ (3.4)
holds true. The three summnads appearing in $\Xi(l)$ are considered as elements in $\operatorname{\mathbb{Z}_{p^{r}}}$. On the other hand, note that the left hand side is known, so $\delta_{l}$ can be computed. Thus, Eq. (3.3) implies that (see [5, Eq. (20)])
$\delta_{l}=H_{0}e_{l}$
in $\operatorname{\mathbb{Z}_{p}}$.
Algorithm 2 Adapted TLN algorithm
Parameters: $H\in\textrm{Mat}_{n-k\times n}(\operatorname{\mathbb{Z}_{p^{r}}})$ full rank parity-check matrix of an $(n,k)$ free linear block code and a decoding algorithm $\partial$ for the linear code given by $H_{0}:=H\ \textrm{mod}(p)$.
Input: A syndrome $s:=H\widehat{v}\in\mathbb{Z}^{n-k}_{p^{r}}$.
Output: $e=\sum_{j=0}^{r-1}e_{j}p^{j}\in\operatorname{\mathbb{Z}^{n}_{p^{r}}}$.
$i=0$
while $i\leq r-1$ do
     $\delta_{i}\leftarrow\dfrac{\Xi(i)}{p^{r-1}}$
     Solve $\delta_{i}=H_{0}e_{i}$ with $\partial$
     $i=i+1$
end while
return $e,v=\widehat{v}-e$
As GV algorithm, TLN algorithm corrects any error pattern $e$ for which each component of its $p$-adic expansion is correctable by $\partial$.
3.4. Rosenthal’s decoding algorithm for convolutional codes over $\operatorname{\mathbb{Z}_{p^{r}}}$
Let $\mathcal{C}\subset\operatorname{\mathbb{Z}_{p^{r}}}[z]^{n}$ be an $(n,k)$ observable convolutional code with PDP and complexity $\delta_{\mathcal{C}}=\delta$, and let $\Sigma=(A,B,C,D)$ be a minimal I/S/O representation over $\operatorname{\mathbb{Z}_{p^{r}}}$. Rosenthal’s decoding algorithm is based on this fact: a code sequence $\left\{c_{t}=\left(\begin{array}[]{c}y_{t}\\ u_{t}\end{array}\right)\right\}$ must satisfy
$\begin{split}&\left(\begin{array}[]{c}y_{\tau+l+1}\\ y_{\tau+l+2}\\ \vdots\\ y_{\tau+l+\Theta}\end{array}\right)-\left(\begin{array}[]{ccccc}D&0&\ldots&&0\\ CB&D&\ddots&&\vdots\\ CAB&CB&\ddots&\ddots&\\ \vdots&&\ddots&\ddots&0\\ CA^{\Theta-2}B&CA^{\Theta-3}B&\ldots&CB&D\end{array}\right)\left(\begin{array}[]{c}u_{\tau+l+1}\\ u_{\tau+l+2}\\ \vdots\\ u_{\tau+l+\Theta}\end{array}\right)=\\ =&\left(\begin{array}[]{c}C\\ CA\\ \vdots\\ CA^{\Theta-1}\end{array}\right)x_{\tau+l+1}\end{split}$ (3.5)
and
$x_{\tau+l+1}-A^{l+1}x_{\tau}=\left(A^{l}B,A^{l-1}B,\ldots,B\right)\left(\begin{array}[]{c}u_{\tau}\\ u_{\tau+1}\\ \vdots\\ u_{\tau+l}\end{array}\right)$ (3.6)
for every $\tau,l,\Theta\in\mathbb{N}$. We will explain now the basis of Rosenthal’s decoding algorithm. Let us make the next assumptions for the moment.
  1. (1)
    $\Phi_{l+1}(\Sigma)$ is surjective and the code $\textrm{Ker}(\Phi_{l}(\Sigma))$ has distance $d_{1}$.
  2. (2)
    $\Psi_{\Theta}(\Sigma)$ is injective and the code $\textrm{Im}(\Psi_{l}(\Sigma))$ has distance $d_{2}$.
  3. (3)
    We have computed the codewords from $t=0$ to $t=\tau-1$ and we know $x_{\tau}$.
  4. (4)
    The inputs ($u^{\prime}s$) in Eq. (3.5) has been received correctly.
  5. (5)
    The vector of outputs ($y$’s) in Eq. (3.6) has weight less than or equal to $[\dfrac{d_{2}-1}{2}]$.
  6. (6)
    The vector of inputs in Eq. (3.6) has weight less than or equal to $[\dfrac{d_{1}-1}{2}]$.
Note that under above assumptions 2), 4) and 5), we may decode $\Psi_{\Theta}(\Sigma)$ (see (3.1)) and find
$\left(\begin{array}[]{c}y_{\tau+l+1}\\ y_{\tau+l+2}\\ \vdots\\ y_{\tau+l+\Theta}\end{array}\right)\textrm{ and }x_{\tau+l+1}$ (3.7)
Now, with $x_{\tau+l+1}$ we can compute the syndrome (see Eq. (3.6))
$\left(A^{l}B,A^{l-1}B,\ldots,B\right)\left(\begin{array}[]{c}\widehat{u}_{\tau}-u_{\tau}\\ \widehat{u}_{\tau+1}-u_{\tau+1}\\ \vdots\\ \widehat{u}_{\tau+1}-u_{\tau+l}\end{array}\right)$
and, under the assumptions 1) and 6), we can decode $\Phi_{l+1}(\Sigma)$ and find
$\left(\begin{array}[]{c}u_{\tau}\\ u_{\tau+1}\\ \vdots\\ u_{\tau+l}\end{array}\right)$
since
$\left(\begin{array}[]{c}\widehat{u}_{\tau}-u_{\tau}\\ \widehat{u}_{\tau+1}-u_{\tau+1}\\ \vdots\\ \widehat{u}_{\tau+1}-u_{\tau+l}\end{array}\right)$
is the unique coset leader in its coset (see 3.1). So far, we have correctly computed the input vectors ($u$’s) from $t=\tau$ to $t=\tau+l+\Theta$ and the output vectors ($y$’s) from $t=\tau+l+1$ to $\tau+l+\Theta$, and we can use the equations (2.3) to compute the remaining outputs from $t=\tau$ to $t=\tau+l$ so the decoding is complete.
In general, above assumptions 5) and 6) might not hold so the decoding process may give no output or may give an error. However, under certain mild assumptions the algorithm ensures that we get the correct output. In fact, [23, Theorem 3.4] can be stated exactly in the same terms for observable convolutional codes with the PDP over $\operatorname{\mathbb{Z}_{p^{r}}}$.
Theorem 3.8.
Let $\mathcal{C}\subset\operatorname{\mathbb{Z}_{p^{r}}[z]}^{n}$ be an $(n,k)$ observable PDP convolutional code, and $\Sigma=(A,B,C,D)$ a minimal I/S/O representation of $\mathcal{C}$. Let $T>\Theta$ be natural numbers such that
  1. (1)
    $A$ is invertible.
  2. (2)
    $\Phi_{T}(A,B)$ is surjective.
  3. (3)
    $\Psi_{\Theta}(A,C)$ is injective.
Then, if we receive a message
$\left\{\left(\begin{array}[]{c}\widehat{y}_{t}\\ \widehat{u}_{t}\end{array}\right)\right\}_{t\geq 0}=\left\{\left(\begin{array}[]{c}y_{t}+f_{t}\\ u_{t}+e_{t}\end{array}\right)\right\}_{t\geq 0}$
such that, for any $\tau\geq 0$, the weight of the errors in the time-window $[\tau,\tau+T-1]$ is bounded by
$\lambda:=\textrm{min}\left(\Big{\lfloor}\dfrac{d_{1}-1}{2}\Big{\rfloor},\Big{\lfloor}\dfrac{T}{2\Theta}\Big{\rfloor}\right),$ (3.9)
it is possible to uniquely compute the transmitted sequence.
Proof.
The proof follows exactly the same steps as in [23], that is, we decode iteratively as stated above for $l=T-\Theta$ (see Algorithm 3). If it is not possible to decode in Eq. (3.5), or the we find that the weight of the computed errors from $t=\tau$ to $t=\tau+T-\Theta$ is larger than $\lambda$, we repeat again the procedure with $l=T-2\Theta$ and substitute $\lambda$ by $\lambda-1$. We proceed in this way until we can decode successfully. As in the field case we can not be sure that the calculation of $x_{\tau+l}$ (for $l=T-h\Theta$ if we needed $h$ iterations) is correct. However, [23, Lemma 3.3] also holds in the ring case, so the conclusion that the state $x_{\tau+\Theta}$ is correct holds as well in our situation. The proof of [23, Lemma 3.3] in the ring case is the same as in the field case. Note that, in our situation, [23, Eq. (3.6)] implies that the distance between the vector of true information words modulo $p$ and the vector of computed information words modulo $p$ is at least $d_{1}$ because the corresponding linear block code is free so its distance remains the same after restricting modulo $p$. Since the distance between vectors is at least the distance between the vectors restricted modulo $p$, we are done because the rest of the argument remains valid in our case. ∎
Algorithm 3 Combined algorithm
Parameters: Minimal I/S/O $\Sigma=(A,B,C,D)$ of an observable convolutional code with PDP, natural numbers $T>\Theta$ satisfying assumptions of Theorem 3.8, $\lambda$ as in Eq. (3.9).
Input:
  1. (1)
    A natural number $h$ with $T\geq h\Theta$.
  2. (2)
    A block of length $T$ of a message $\left\{\left(\begin{array}[]{c}\widehat{y}_{t}\\ \widehat{u}_{t}\end{array}\right)\right\}_{t\geq 0}=\left\{\left(\begin{array}[]{c}y_{t}+f_{t}\\ u_{t}+e_{t}\end{array}\right)\right\}_{t\geq 0}$ (let us say, from $t=\tau$ to $t=\tau+T-1$)
Output:
  1. (1)
    An estimation of the first $\Theta$ words $\left\{\left(\begin{array}[]{c}\tilde{y}_{t}\\ \tilde{u}_{t}\end{array}\right)\right\}_{t=\tau}^{\tau+\Theta-1}$ and the state $x_{\tau+\Theta}$ if possible,
  2. (2)
    or an ERROR message together with $h\leftarrow h+1$.
$\lambda\leftarrow\lambda-h$
$\Theta\leftarrow h\Theta$
$l\leftarrow T-h\Theta$
if Decoding on (3.5) is successful then
     Compute syndrome (3.6)
     if Syndrome decoding on (3.5) is successful then
         Use (2.3) to compute the remaining codewords
     else
         return (ERROR, $h\leftarrow h+1$)
     end if
else
     return (ERROR, $h\leftarrow h+1$)
end if
if $\sum_{\tau}^{\tau+T-\Theta}(\omega_{H}(f_{t})+\omega_{H}(e_{t}))>\lambda$ then
     return (ERROR, $h\leftarrow h+1$)
else
     return ( $\left\{\left(\begin{array}[]{c}\tilde{y}_{t}\\ \tilde{u}_{t}\end{array}\right)\right\}_{t=\tau}^{\tau+\Theta-1}$, state $x_{\tau+\Theta}$ )
end if
Corollary 3.10.
Let $\mathcal{C}\subset\operatorname{\mathbb{Z}_{p^{r}}[z]}^{n}$ be an $(n,k)$ observable PDP convolutional code, and $\Sigma=(A,B,C,D)$ a minimal I/S/O representation of $\mathcal{C}$. Let $T>\Theta$ be natural numbers satisfying the same conditions as in Theorem 3.8. Let $\partial_{gen}$, $\partial_{par}$ be decoding algorithms for the linear codes (over $\operatorname{\mathbb{Z}_{p}}$) determined by $\textrm{Im}(\Psi_{\Theta,0})$ (the first coefficient in the $p$-adic expansion of $\Psi_{\Theta}(A,C)$), and $\textrm{Ker}(\Phi_{T,0})$ (the first coefficient of the $p$-adic expansion of $\Phi_{T}(A,B)$) and let $t_{gen},t_{par}$ be the correction capabilities of $\partial_{gen}$, $\partial_{par}$, respectively. Then, if we receive a message
$\left\{\left(\begin{array}[]{c}\widehat{y}_{t}\\ \widehat{u}_{t}\end{array}\right)\right\}_{t\geq 0}=\left\{\left(\begin{array}[]{c}y_{t}+f_{t}\\ u_{t}+e_{t}\end{array}\right)\right\}_{t\geq 0}$
such that, for any $\tau\geq 0$,
  1. (1)
    the weight of the errors $\left\{\left(\begin{array}[]{c}f_{t}\\ e_{t}\end{array}\right)\right\}_{t=\tau}^{\tau+T-1}$ is bounded by $\lambda:=\textrm{min}\left(\Big{\lfloor}\dfrac{d_{1}-1}{2}\Big{\rfloor},\Big{\lfloor}\dfrac{T}{2\Theta}\Big{\rfloor}\right)$.
  2. (2)
    the sequence of the $i$th coefficients in the $p$-adic expansion of the vectors of the error sequence $(f_{\tau},\ldots,f_{\tau+T-1})$ has weight less or equal to $t_{gen}$,
  3. (3)
    the sequence of the $i$th coefficients in the $p$-adic expansion of the vectors of the error sequence $(e_{\tau},\ldots,e_{\tau+T-1})$ has weight less or equal to $t_{par}$,
it is possible to uniquely compute the transmitted sequence.
Proof.
It follows from Theorem 3.8 and applying Algorithm 1 and Algorithm 2 in the the decoding steps of Algorithm 3. ∎
4. Performance of the algorithm
Let $\mathcal{C}\subset\operatorname{\mathbb{Z}_{p^{r}}[z]}^{n}$ be an $(n,k)$ observable PDP convolutional code, and $\Sigma=(A,B,C,D)$ a minimal I/S/O representation of $\mathcal{C}$. Let $T>\Theta$ be natural numbers as in Theorem 3.8.
Regarding the performance of the algorithm, it is important to note that, essentially, it depends on the theoretical capacity of the convolutional code modulo $p$. More precisely, in case
$\lambda=\Big{\lfloor}\dfrac{d_{1}-1}{2}\Big{\rfloor},$
the theoretical performance of the algorithm depends on the minimum distance of the linear code whose parity-check matrix is given by $\Psi_{\Theta}(A,C)$. Since this code is free, its minimum distance coincides with with the linear code over $\operatorname{\mathbb{Z}_{p}}$ whose parity-check matrix is given by $\Psi_{\Theta}(A,C)\textrm{mod}\ p$. This, together with [17, Remark 7], allow to conclude (as in [23, Remark 3.6]) that decoding is possible if no more than
$\dfrac{d_{\textrm{free}}(\mathcal{C}\ \textrm{mod}\ p)-1}{2}$
erors occur in any time interval of length $T$.
5. Conclusion
In this article, Rosenthal’s algorithm for the decoding process of a convolutional code using I/S/O representations has been generalized to $\operatorname{\mathbb{Z}_{p^{r}}}$. As future work, the goal is to extend the use of I/S/O representations for decoding over erasure channels to the context of convolutional codes over rings.